The lengths of a professor's classes has a continuous uniform distribution between 50.0 min and 52.0 min. If one such class is randomly selected, find the probability that the class length is less than 51.8 min.

P(X < 51.8) =
(Report answer accurate to 2 decimal places.)
Answer #1

Solution :

Given that,

a = 50

b = 52

P(x < c) = (c - a) / (b - a)

P(x < 51.8) = (51.8 - 50) / (52 - 50) =0.90

P(X < 51.8) =0.90